OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –6 3 3 6 Line 2 x y –3 1 3 3 A. 0 B. exactly 1 C. exactly 2 D. infinitely many for line 1 i got 3/9 and for line 2 i got 2/0

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

help plz

Directrix (directrix):

I can check your answers. Hold on.

OpenStudy (anonymous):

ok

Directrix (directrix):

Is this >>for line 1 i got 3/9 and for line 2 i got 2/0 >> Are those slopes or what?

OpenStudy (anonymous):

yea they r

Directrix (directrix):

The equation of the first line is this: y = (1/3) * x + 5 I got this using the online calculator that gives the line equation when two points are put into the calculator. It is at this URL: http://www.mathportal.org/calculators/analytic-geometry/two-point-form-calculator.php

Directrix (directrix):

Now, for the next line's equation.

Directrix (directrix):

Line Two's equation is this: y = (1/3) * x + 2

Directrix (directrix):

Look at the two lines together. y = (1/3) * x + 5 y = (1/3) * x + 2

OpenStudy (anonymous):

i think it is D

Directrix (directrix):

They have the same slope but different y-intercepts. They are parallel lines. Recall that two lines with the same slope are parallel or the same line. These cannot be the same line because they have different y-intercepts.

OpenStudy (anonymous):

so C

Directrix (directrix):

I don't see that it is D. Hold on and let's look at the graphs of the two lines on the same axes.

Directrix (directrix):

Not C.

Directrix (directrix):

Click on this link and look at the blue and red lines. How many points do they have in common? http://www.wolframalpha.com/input/?i=+y+%3D+%281%2F3%29+*+x+%2B+5%2C+y+%3D+%281%2F3%29+*+x+%2B+2

Directrix (directrix):

I have attached the graphs here.

OpenStudy (anonymous):

A

Directrix (directrix):

How many solutions does this system of equations have? ZERO

OpenStudy (anonymous):

yup thanks

Directrix (directrix):

I agree with option A.

Directrix (directrix):

You are welcome.